# Economic DAO Governance: A Contestable Control Approach

### v3

@@ A Contestable Control Approach
Jeff Strnad [^1] Charles A. Beardsley Professor of Law, Stanford University. Correspondence email: jeffstrnad5@gmail.com. I am grateful for valuable comments from Bobby Bartlett, Dan Boneh, Albert Choi, Silke Elrifai, Mike Guttentag, Jake Hartnell, Henry Hu, Mark Lemley, Sven Riva, Roberta Romano, Martin Schmidt, Mike Simkovic, Connor Spelliscy, and Jay Yu, and from participants at the DAO @ LexTech Institute 2024 Conference, the 2024 Stanford Blockchain Governance Summit, the Stanford Law School Faculty Workshop, the TLDR Conference 2024, and the 2024 American Law and Economics Association Annual Meeting. I have also benefited greatly from conversations with Andy Hall.
In this article, we propose a new form of DAO governance that uses a sequential auction mechanism to overcome entrenched control issues that have emerged for DAOs by creating a regime of temporary contestable control. The mechanism avoids potential public choice problems inherent in voting approaches but at the same time provides a vehicle that can enhance and secure value that inheres to DAO voting and other DAO non-market governance procedures. It is robust to empty voting and is code feasible. It facilitates the ability of DAOs to meet their normative and operational goals in the face of diverse regulatory approaches. Designed to shift control to the party with the most promising business plan, at the same time it distributes social surplus in a way that tends to promote investment by other parties.
June 9, 2024 Version
@@ 2.1 Addressing Social Choice Problems
Most generally, there is the Gibbard-Satterthwaite Theorem derived by Gibbard (1973) and Satterthwaite (1975) which states that if individual preference orderings are complete and transitive but otherwise unrestricted and there are at least three alternatives, then a direct mechanism is dominant strategy incentive-compatible if and only if it is dictatorial.999 Borgers (2015) provides a concise discussion and proof of this Theorem and its significance. In a rough sense what the Theorem means is that with unrestricted preferences, a mechanism in which each individual votes sincerely, recording their actual preferences, will only work if there is a single individual who decides everything or if there are no more than two alternatives. Once strategic voting enters the picture, results can become unpredictable, difficult to estimate, and possibly very unrepresentative compared to voters’ actual preferences. [^10] The discussion by Tabarrok and Spector (1999) about how strategic voting might have affected the U.S. Presidential election of 1860 under various voting regimes is a good example.
@@ 3 Conceptual Overview
We assume the subject DAO is an economic DAO. We consider a setting similar to Burkart and Lee (2021) in which various parties can use costly effort to increase the value of the DAO. In particular, suppose each party can engage in various projects that consist of expending effort, labor, and resources equivalent to $C$ monetary units, in order to increase the value of the DAO by $N(C)$ monetary units. [^20] These costs are external to the DAO and are borne by the party engaging in the project. If the party uses DAO resources such as DAO treasury assets to execute the project, use of these resources will diminish the value of the DAO directly, reducing $N(C)$ rather than being an addition to $C$. Suppose that the DAO has $q$ tokens outstanding and that the current market price per token is $P_{0}$. Define $V=V(C)=N(C)/q+P_{0}$ to be the token value emerging from a particular business plan. This business plan will create social surplus $\psi=(V-P_{0})q-C$. Consider that expectations about the nature of future value-additive projects and about the distribution of social surplus from those projects will affect the initial funding value of the DAO, possibly being critical to having enough funding to start up at all. In other words, the expected later treatment and facilitation of value-added projects will have an investment impact on initial funding for DAOs, and also, in an obvious way, on the on-going investment value of the DAO, which continues to depend on possible future projects.
@@
The mechanism created here addresses the potential value-reducing free-riding of both types through two devices. First, by creating a freeze-out feature as part of an auction, the mechanism allows the project creator to buy a proportion of the other token holders’ positions for no premium. Those token holders will earn the full amount of the expected total added token value on the retained proportion that is not purchased by the creator. This feature makes the auction equivalent to a merger in which token holders receive only a portion of the expected total added token value on their token position shares because of the terms of the merger which dictate the proportional stake in the merged entity that they receive in exchange for their shares in one of the precursor entities. [^24] Although the mechanism creates an outcome somewhat analogous to a merger, no actual merger is involved. Trading in the DAO remains continuous, and the life of the DAO goes forward. A merger in which the acquiring party buys all of the DAO shares would mean the end of the DAO in its current form. It might be that the acquiring party, which may be a shell, is itself set up to be a DAO, perhaps one with a new set of smart contracts meant to upgrade the acquired DAO. We leave exploration of this possibility to future work, including the case in which the selling token holders receive tokens in the new DAO instead of a cash-equivalent in exchange for the tokens in the old DAO. The bidding mechanism creates an incentive for bidders to limit the amount of total added token value that they attempt to claim via the freeze-out feature to the lowest possible value, the amount required to cover their costs taking into account the expected gain on their toehold position. As discussed in the next section, there typically will be some leakage because the winning bidder will have some scope to go beyond that limit.
@@
The approach here flows from a very general deposit-based model: There is a single auction deposit, $D_{a}$, and a deposit forfeit function $\Phi(D_{a},P_{0},S,X)$ where $P_{0}$ is the token price when the auction is initiated, $S$ is the value that the winning bidder claims is attainable, and $X$ is an outcome, typically the token price when the winning bidder’s control terminates. The first three parameters in the function are set by the time the auction ends and the winning bidder takes control. The deposit forfeit function specifies the amount of the deposit that is forfeited for each possible value of $X$ given the three auction parameters. The function is not required to be continuous in $X$, and we will see that discontinuities can play an important role in creating incentives for control parties.
@@ 4.1.1 The Basic Auction Mechanism
A basic auction is initiated at a time $T_{0}$ by a first bidder making a bid. The basic auction is open for the competitive bidding process up until a winner is determined at some later time $T_{1}$, with the fixed total time length $T_{auction}=T_{1}-T_{0}$ of the bidding period specified in advance by the DAO Code. Suppose that at $T_{0}$ the price of the DAO governance token is $P_{0}$ and $q$ tokens are outstanding. A bid, $\beta=(S,R,t_{b})$, consists of three parameters: a value claim, $S\geq P_{0}$, per token; $t_{b}$, the proportion of the q tokens held by the bidder at $T_{0}$, hereinafter termed the bidder’s toehold; and a surplus claim, $R\leq(t_{m}-t_{b})(S-P_{0})$ where $t_{m}$ is a parameter set by the DAO Code. The bidder’s toehold, $t_{b}q$, is deposited into the applicable smart contract simultaneously with the bid, becoming part of the bidder’s required token deposit under the mechanism. The basic auction is an English auction, ascending in the auction parameter $A=(1-t_{b})(S-P_{0})q-R$. All prices and quantities such as $S,~{}P_{0},~{}\mbox{and}~{}R$ are denominated in the units of a particular reference fiat currency designated by the DAO Code.
@@
Some aspects of the standard case generalize. Under the mechanism, the value deposit forfeit function is always set in such a manner that the token holders other than the bidder will receive at least the full gain $S-P_{0}$ on their $(1-t_{d})$ share of the tokens whether or not the business plan is successfully implemented. If the business plan succeeds, the other token holders realize the full gain because the market price increases to at least $S$, fulfilling the value claim. If not, operation of the value deposit forfeit function ensures that any deficit is more than made up from the value deposit. [^33] In cases in which the bidder’s control ends in a subsequent auction, implementation of any refund may be delayed until after subsequent control periods and may become moot in whole or in part due to the performance of the token price under the guidance of the subsequent control parties. To the extent the refund is not paid out as part of the transition to the next control period, the mechanism preserves it as an obligation. See subsections 4.2 and 4.3 infra. Thus, the value deposit combined with the operation of the value deposit forfeit function transforms the bidder’s value claim into a commitment to deliver the promised increase in token value to the other token holders. The fact that this deposit is held by the applicable smart contract makes this commitment credible and immediately enforceable in a code feasible manner.
@@
Consistent with the lack of centralized management in a DAO, the auction is designed to be self-executing through the Auction Contracts. To reach that goal, the auction mechanism must be code feasible. All four deposits are useful in that respect. Because of the purchase deposit requirement, non-payment cannot derail the auction. There is no need to have recourse outside of the Code to legal process for purposes of collection. As noted in the margin, [^37] See note 32 supra, which states the equivalent derivatives position for a case in which that position is a simple options spread. In general, there will be an equivalent derivatives position, but it may not be simple. the value deposit, combined with the value deposit forfeit function, serves a function that also could be accomplished through derivatives. Use of a deposit strategy eliminates the need for counterparties and the possible need to enforce counterparty compliance with the option contracts, elements that may not be code feasible without a great deal of added complexity or at all. Similarly, the token and surety deposits substitute for mechanisms that would rely on conventional derivative contracts and escrow arrangements enforced through the legal system.
@@
    The DAO Code creates a dynamic vote pool consisting of $v_{T}(q_{T},t_{d}q)$ additional voting rights assigned to the bidder, an amount that is adjusted continuously based on $q_{T}$, the number of tokens outstanding and eligible to vote at each future time $T$, to ensure that the bidder retains majority control of the DAO. $v_{T}=(0.5q_{T}+1)-t_{d}q$ where $q$ is the number of tokens outstanding and eligible to vote as of time $T_{1}$ when the auction ends. The votes in the pool are empty votes because they are not matched with the corresponding economic interest inherent in the tokens. [^39] The choice of giving the winning bidder majority control (>.$>.5$ of the votes) implicit in this arrangement leaves open the possibility that the winning bidder will not be able to prevail on DAO issues that require a threshold greater than 50% to pass the applicable proposal. The exact percentage established by the DAO Code for the vote pool needs to be sensitive to the hierarchy of voting thresholds and the associated issues set for the DAO more generally. If a less conventional choice method such as quadratic voting is employed by the DAO, the dynamic voting pool must grant whatever number of additional voting rights is required to establish control.
@@ 4.1.2 The Optimal Business Plan and Bidding Strategy
Suppose a potential bidder can implement a business plan $(V,C)$ that involves expending effort, consisting of labor and resources equivalent to $C$ monetary units, that will result in a token value $V>P_{0}$. This business plan creates social surplus $\psi(V,C)=(V-P_{0})q-C$. Suppose that among all of the potential bidder’s possible business plans, the business plan $(V^{*},C^{*})$ creates the largest amount of social surplus, $\psi^{*}=(V^{*}-P_{0})q-C^{*}$.404040 We assume that the auction takes place in a “private values” setting with asymmetric information. Each bidder’s potential set of business plans and their value-enhancing potential are known only to the bidder. All bidders agree on the value of the current operation. If there is a common values element in which bidders have different signals concerning the value of the current operation and bid based on perceived undervaluation at price $P_{0}$, then the efficiency and surplus-distribution characteristics of the auction mechanism may be affected. However, choosing the English auction as a mechanism tends to make any such impacts benign or even beneficial. See note 66 infra.
@@
In Figure 1, the business plan $(V,C)$ induces a value claim $S=V$ and a surplus claim value of $R$ just big enough to cover the cost $C$, which results in the winning bidder purchasing $t_{f}q$ tokens at $P_{0}$ from the $T_{1}$ token holders through the freeze-out mechanism. This purchase results in the bidder realizing added token value equal to the blue $R$ rectangle in the figure. The green rectangle $\Pi_{t_{f}}$ is the added token value that the bidder realizes from the toehold position $t_{b}q$. The bidder’s total added token value is the sum of the green and blue rectangles, which is equal in this instance to the red rectangle representing the cost $C$. The area $F$ in the figure represents the minimum added token value that must be granted to the other post-$T_{1}$ token holders, the proportion $1-t_{m}$ of the total, and these free-riding shareholders also realize added token values equal to the unlabeled white rectangle between the $R$ rectangle and the $F$ rectangle.
@@
$A^{*}$ proxies for bidder types, and a bidder with a lower $A^{*}$ cannot profitably mimic one with a higher $A^{*}$. For bidder $i$, the strongest possible bid, $A_{i}^{*}$, is associated with zero profits. A higher bid results in net losses. As the English auction unfolds, if each bidder $i$ moves up until the level $A_{i}^{*}$ and then drops out, $V_{i}^{*}$ and $C_{i}^{*}$ will be evident because the bid also reveals $t_{b}$. For the winning bidder, only $V_{1}^{*}$ and an upper bound on $C_{1}^{*}$ will be evident because it is possible that the winning bid, $A_{1}<A_{1}^{*}$ so that $\Pi_{t_{b}}+R_{1}>C_{1}^{*}$. In that case, the winning bidder is receiving more added token value than is necessary to induce that bidder to implement the best possible business plan. This situation appears to fall short of the first best which requires that all of the social surplus remains with the free riding token holders in order to optimize initial and on-going investment in the DAO. But, as discussed in subsection 5.1, the full picture is more complex.

### v4

# +1.5in-1.2in

@@ A Contestable Control Approach
Jeff Strnad***Charles A. Beardsley Professor of Law, Stanford University. Correspondence email: jeffstrnad5@gmail.com. I am grateful for valuable comments from Bobby Bartlett, Dan Boneh, Albert Choi, Silke Elrifai, Mike Guttentag, Jake Hartnell, Henry Hu, Takuma Iwasaki, Mark Lemley, Sven Riva, Roberta Romano, Martin Schmidt, Mike Simkovic, Connor Spelliscy, and Jay Yu, and from participants at the DAO @ LexTech Institute 2024 Conference, the 2024 Stanford Blockchain Governance Summit, the Stanford Law School Faculty Workshop, the TLDR Conference 2024, the 2024 American Law and Economics Association Annual Meeting, Edge Esmeralda 2024, UBRI Connect 2024, and the Hackers Congress Paralelní Polis 2024. I have also benefited greatly from conversations with Andy Hall.
In this article, we propose a new form of DAO governance that uses a sequential auction mechanism to overcome entrenched control issues that have emerged for DAOs by creating a regime of temporary contestable control. The mechanism avoids potential public choice problems inherent in voting approaches but at the same time provides a vehicle that can enhance and secure value that inheres to DAO voting and other DAO non-market governance procedures. It is robust to empty voting and is code feasible. It not only facilitates the ability of DAOs to meet their normative and operational goals in the face of diverse regulatory approaches, but also strengthens the case for creating a less burdensome but at least equally effective regulatory regime for DAOs that employ the mechanism. Designed to shift control to the party with the most promising business plan, at the same time it deters value destruction by control parties, maximizes social surplus, and distributes that surplus in a way that tends to promote investment by other parties both at start up and on an on-going basis.
Keywords: blockchain governance, decentralized autonomous organizations, DAO, DAO governance, control auctions, cryptocurrency regulation, DAO regulation, auction mechanisms, empty voting
November 25, 2024
@@ 2.1 Addressing Social Choice Problems
Most generally, there is the Gibbard-Satterthwaite Theorem derived by Gibbard (1973) and Satterthwaite (1975) which states that if individual preference orderings are complete and transitive but otherwise unrestricted and there are at least three alternatives, then a direct mechanism is dominant strategy incentive-compatible if and only if it is dictatorial.888 Borgers (2015) provides a concise discussion and proof of this Theorem and its significance. In a rough sense what the Theorem means is that with unrestricted preferences, a mechanism in which each individual votes sincerely, recording their actual preferences, will only work if there is a single individual who decides everything or if there are no more than two alternatives. Once strategic voting enters the picture, results can become unpredictable, difficult to estimate, and possibly very unrepresentative compared to voters’ actual preferences. [^9] The discussion by Tabarrok and Spector (1999) about how strategic voting might have affected the U.S. Presidential election of 1860 under various voting regimes is a good example.
@@ 3 Conceptual Overview
We assume the subject DAO is an economic DAO. We consider a setting similar to Burkart and Lee (2021) in which various parties can use costly effort to increase the value of the DAO. In particular, suppose each party can engage in various projects that consist of expending effort, labor, and resources equivalent to $C$ monetary units, in order to increase the value of the DAO by $N(C)$ monetary units. [^19] These costs are external to the DAO and are borne by the party engaging in the project. If the party uses DAO resources such as DAO treasury assets to execute the project, use of these resources will diminish the value of the DAO directly, reducing $N(C)$ rather than being an addition to $C$. Suppose that the DAO has $q$ tokens outstanding and that the current market price per token is $P_{0}$. Define $V=V(C)=N(C)/q+P_{0}$ to be the token value emerging from a particular business plan. This business plan will create social surplus $\psi=N(C)-C=(V-P_{0})q-C$. Consider that expectations about the nature of future value-additive projects and about the distribution of social surplus from those projects will affect the initial funding value of the DAO, possibly being critical to having enough funding to start up at all. In other words, the expected later treatment and facilitation of value-added projects will have an investment impact on initial funding for DAOs, and also, in an obvious way, on the on-going investment value of the DAO, which continues to depend on possible future projects.
@@
The mechanism created here addresses the potential value-reducing free-riding of both types through two devices. First, by creating a freeze-out feature as part of an auction, the mechanism addresses the Grossman-Hart free-rider problem by allowing the project creator to buy a proportion of the other token holders’ positions for no premium. Those token holders will earn the full amount of the expected total added token value on the retained proportion that is not purchased by the creator. This feature makes the auction equivalent to a merger in which token holders receive only a portion of the expected total added token value on their token position shares because of the terms of the merger which dictate the proportional stake in the merged entity that they receive in exchange for their shares in one of the precursor entities. [^23] Although the mechanism creates an outcome somewhat analogous to a merger, no actual merger is involved. Trading in the DAO remains continuous, and the life of the DAO goes forward. A merger in which the acquiring party buys all of the DAO shares would mean the end of the DAO in its current form. It might be that the acquiring party, which may be a shell, is itself set up to be a DAO, perhaps one with a new set of smart contracts meant to upgrade the acquired DAO. We leave exploration of this possibility to future work, including the case in which the selling token holders receive tokens in the new DAO instead of a cash-equivalent in exchange for the tokens in the old DAO. The bidding mechanism creates an incentive for bidders to limit the amount of total added token value that they attempt to claim via the freeze-out feature to the lowest possible value, the amount required to cover their costs taking into account the expected gain on their toehold position. As discussed in the next section, there typically will be some leakage because the winning bidder will have some scope to go beyond that limit.
@@
The approach here flows from a very general deposit-based model: There is a single auction deposit, $D_{a}$, and a deposit forfeit function $\Phi(D_{a},P_{0},S,X)$ where $P_{0}$ is the token price when the auction is initiated, $S$ is the value per token that the winning bidder claims is attainable, and $X$ is an outcome, typically the token price when the winning bidder’s control terminates. The first three parameters in the function are set by the time the auction ends and the winning bidder takes control. The deposit forfeit function specifies the amount of the deposit that is forfeited for each possible value of $X$ given the three auction parameters. The function is not required to be continuous in $X$, and we will see that discontinuities can play an important role in creating incentives for control parties.
@@ 4.1.1 The Basic Auction Mechanism
A basic auction is initiated at a time $T_{0}$ by a first bidder making a bid. The basic auction is open for the competitive bidding process up until a winner is determined at some later time $T_{1}$, with the fixed total time length $T_{auction}=T_{1}-T_{0}$ of the bidding period specified in advance by the DAO Code. Suppose that at $T_{0}$ the price of the DAO governance token is $P_{0}$ and $q$ tokens are outstanding. A bid, $\beta=(S,R,t_{b})$, consists of three parameters: a value claim, $S\geq P_{0}$, per token; $t_{b}$, the proportion of the q tokens held by the bidder at $T_{0}$, hereinafter termed the bidder’s toehold; and a surplus claim, $R\leq(t_{m}-t_{b})(S-P_{0})q$ where $t_{m}$ is a parameter set by the DAO Code that, as discussed in subsection 4.1.6, plays a role in assuring adequate market liquidity. The bidder’s toehold, $t_{b}q$, is deposited into the applicable smart contract simultaneously with the bid, becoming part of the bidder’s required token deposit under the mechanism. The basic auction is an English auction, ascending in the auction parameter $A=(1-t_{b})(S-P_{0})q-R$. All prices and quantities such as $S,~{}P_{0},~{}\mbox{and}~{}R$ are denominated in the units of a particular reference fiat currency designated by the DAO Code.
@@
Some aspects of the standard case generalize. Under the mechanism, the value deposit forfeit function is always set in such a manner that the token holders other than the bidder will receive at least the full gain $S-P_{0}$ per token on their $(1-t_{d})$ share of the tokens whether or not the business plan is successfully implemented. If the business plan succeeds, the other token holders realize the full gain because the market price increases to at least $S$, fulfilling the value claim. If not, operation of the value deposit forfeit function ensures that any deficit is more than made up from the value deposit. [^32] In cases in which the bidder’s control ends in a subsequent auction, implementation of any refund may be delayed until after subsequent control periods and may become moot in whole or in part due to the performance of the token price under the guidance of the subsequent control parties. To the extent the refund is not paid out as part of the transition to the next control period, the mechanism preserves it as an obligation. See subsections 4.2 and 4.3 infra. Thus, the value deposit combined with the operation of the value deposit forfeit function transforms the bidder’s value claim into a commitment to deliver the promised increase in token value to the other token holders. The fact that this deposit is held by the applicable smart contract makes this commitment credible and immediately enforceable in a code feasible manner.
@@
Consistent with the lack of centralized management in a DAO, the auction is designed to be self-executing through the Auction Contracts. To reach that goal, the auction mechanism must be code feasible. All four deposits are useful in that respect. Because of the purchase deposit requirement, non-payment cannot derail the auction. There is no need to have recourse outside of the Code to legal process for purposes of collection. As noted in the margin, [^36] See note 31 supra, which states the equivalent derivatives position for a case in which that position is a simple options spread. In general, there will be an equivalent derivatives position, but it may not be simple. the value deposit, combined with the value deposit forfeit function, serves a function that also could be accomplished through derivatives. Use of a deposit strategy eliminates the need for counterparties and the possible need to enforce counterparty compliance with the option contracts, elements that may not be code feasible without a great deal of added complexity or at all. Similarly, the token and surety deposits substitute for mechanisms that would rely on conventional derivative contracts and escrow arrangements enforced through the legal system.
@@
    The DAO Code creates a dynamic vote pool consisting of $v_{T}(q_{T},t_{d}q)$ additional voting rights assigned to the bidder, an amount that is adjusted continuously based on $q_{T}$, the number of tokens outstanding and eligible to vote at each future time $T$, to ensure that the bidder retains majority control of the DAO. The required condition is $v_{T}+t_{d}q>0.5(q_{T}+v_{T})$, which can be met by setting $v_{T}=q_{T}-2t_{d}q+1$ where $q$ is the number of tokens outstanding and eligible to vote as of time $T_{1}$ when the auction ends. The votes in the pool are empty votes because they are not matched with the corresponding economic interest inherent in the tokens. [^38] The choice of giving the winning bidder majority control (>.$>.5$ of the votes) implicit in this arrangement leaves open the possibility that the winning bidder will not be able to prevail on DAO issues that require a threshold greater than 50% to pass the applicable proposal. The exact percentage established by the DAO Code for the vote pool needs to be sensitive to the hierarchy of voting thresholds and the associated issues set for the DAO more generally. If a less conventional choice method such as quadratic voting is employed by the DAO, the dynamic voting pool must grant whatever number of additional voting rights is required to establish control.
@@ 4.1.2 The Optimal Business Plan and Bidding Strategy
Suppose a potential bidder can implement a business plan $(V,C)$ that involves expending effort, consisting of labor and resources equivalent to $C$ monetary units, that will result in a token value $V>P_{0}$. This business plan creates social surplus $\psi(V,C)=(V-P_{0})q-C$. Suppose that among all of the potential bidder’s possible business plans, the business plan $(V^{*},C^{*})$ creates the largest amount of social surplus, $\psi^{*}=(V^{*}-P_{0})q-C^{*}$.393939 We assume that the auction takes place in a “private values” setting with asymmetric information. Each bidder’s potential set of business plans and their value-enhancing potential are known only to the bidder. All bidders agree on the value of the current operation. If there is a common values element in which bidders have different signals concerning the value of the current operation and bid based on perceived undervaluation at price $P_{0}$, then the efficiency and surplus-distribution characteristics of the auction mechanism may be affected. However, choosing the English auction as a mechanism tends to make any such impacts benign or even beneficial. See note 65 infra.
@@
In Figure 1, the business plan $(V,C)$ induces a value claim $S=V$ and a surplus claim value of $R$ just big enough to cover the cost $C$, which results in the winning bidder purchasing $t_{f}q$ tokens at $P_{0}$ from the $T_{1}$ token holders through the freeze-out mechanism. This purchase results in the bidder realizing added token value equal to the blue $R$ rectangle in the figure. The green rectangle $\Pi_{t_{b}}$ is the added token value that the bidder realizes from the toehold position $t_{b}q$. The bidder’s total added token value is the sum of the green and blue rectangles, which is equal in this instance to the red rectangle representing the cost $C$. The area $F$ in the figure represents the minimum added token value that must be granted to the other post-$T_{1}$ token holders, the proportion $1-t_{m}$ of the total, and these free-riding token holders also realize added token value equal to the unlabeled white rectangle between the $R$ rectangle and the $F$ rectangle.
@@
$A^{*}$ proxies for bidder types, and a bidder with a lower $A^{*}$ cannot profitably mimic one with a higher $A^{*}$. For bidder $i$, the strongest possible bid, $A_{i}^{*}$, is associated with zero profits. A higher bid results in net losses. As the English auction unfolds, if each bidder $i$ moves up until the level $A_{i}^{*}$ and then drops out, $V_{i}^{*}$ and $C_{i}^{*}$ will be evident because the bid also reveals $t_{b}$. [^40] As discussed in subsection 4.1.6, it is possible that the second highest bidder submitting a bid with $A_{2}>A_{2}^{*}$ by understating the surplus claim required to break even is a profitable strategy. Although the bidder risks an operational loss if the bidder wins the auction, the bidder will retain a larger proportion of its toehold and earn the ensuing profit per token, $V-P_{0}$, on that larger position if it loses. The winning bidder faces a higher floor, $A_{2}>A_{2}^{*}$, instead of $A_{2}^{*}$ and cannot claim as much of the available social surplus in that instance. For the winning bidder, only $V_{1}^{*}$ and an upper bound on $C_{1}^{*}$ will be evident because it is possible that the winning bid, $A_{1}<A_{1}^{*}$ so that $\Pi_{t_{b}}+R_{1}>C_{1}^{*}$. In that case, the winning bidder is receiving more added token value than is necessary to induce that bidder to implement the best possible business plan. This situation appears to fall short of the first best which requires that all of the social surplus remains with the free riding token holders in order to optimize initial and on-going investment in the DAO. But, as discussed in subsection 5.1, the full picture is more complex.